How do I solve this? (please include how calculations were made and account titles, and the numbers that go with those accounts) I am very confused.

Exercise 10-24

On December 31, 2014, Travis Tritt Inc. has a machine with a book value of $1,406,240. The original cost and related accumulated depreciation at this date are as follows.

Machine
$1,944,800
Less: Accumulated depreciation
538,560
Book value
$1,406,240


Depreciation is computed at $89,760 per year on a straight-line basis.

Presented below is a set of independent situations. For each independent situation, indicate the journal entry to be made to record the transaction. Make sure that depreciation entries are made to update the book value of the machine prior to its disposal.
A. fire completely destroys the machine on August 31, 2015. An insurance settlement of $643,280 was received for this casualty. Assume the settlement was received immediately.
B. On April 1, 2015, Tritt sold the machine for $1,555,840 to Dwight Yoakam Company.

C.On July 31, 2015, the company donated this machine to the Mountain King City Council. The fair value of the machine at the time of the donation was estimated to be $1,645,600.



























I do not
Answer #1

A. fire completely destroys the machine on August 31, 2015. An insurance settlement of $643,280 was received for this casualty. Assume the settlement was received immediately.

Depreciation Expenses till August 31 = 89760*8/12 = $ 59,840

Date Account Title & Explaination Debit Credit
August 31, 2015 Depreciation Expenses 59840
Accumulated Depreciation 59840

Accumulated Depreciation after this entry = 538560 + 59840 = $ 598400

Date Account Title & Explaination Debit Credit
August 31, 2015 Accumulated Depreciation 598400
Cash 643280
Loss on destroy due to fire 703120
machine 1944800


B. On April 1, 2015, Tritt sold the machine for $1,555,840 to Dwight Yoakam Company.

Depreciation Expenses till April 1 = 89760*3/12 = $ 22440

Date Account Title & Explaination Debit Credit
April 1, 2015 Depreciation Expenses 22440
Accumulated Depreciation 22440

Accumulated Depreciation after this entry = 538560 + 22440 = $ 561000

Date Account Title & Explaination Debit Credit
April 1, 2015 Accumulated Depreciation 561000
Cash 1555840
Gain on Sale of Machine 172040
machine 1944800


C.On July 31, 2015, the company donated this machine to the Mountain King City Council. The fair value of the machine at the time of the donation was estimated to be $1,645,600.

Depreciation Expenses till   July 31, 2015 = 89760*7/12 = $ 52360

Date Account Title & Explaination Debit Credit
July 31, 2015 Depreciation Expenses 52360
Accumulated Depreciation 52360

Accumulated Depreciation after this entry = 538560 + 52360 = $ 590920

Date Account Title & Explaination Debit Credit
July 31, 2015 Accumulated Depreciation 590920
Donation 1353880
Machine 1944800